It is widely accepted that people are a little taller in the morning than at night. Here we perform a test on how big the difference is. In a sample of 30 adults, the morning height and evening height are given in millimeters (mm) in the table below. Use this data to test the claim that, on average, people are more than 10 mm taller in the morning than at night. Test this claim at the 0.01significance level.



(a) The claim is that the mean difference (x - y) is more than 10 mm (μd > 10). What type of test is this?

This is a two-tailed test.This is a left-tailed test.     This is a right-tailed test.


(b) What is the test statistic? Round your answer to 2 decimal places.
t

d

=  

(c) Use software to get the P-value of the test statistic. Round to 4 decimal places.
P-value =  

(d) What is the conclusion regarding the null hypothesis?

reject H0fail to reject H0     


(e) Choose the appropriate concluding statement.

The data supports the claim that, on average, people are more than 10 mm taller in the morning than at night.There is not enough data to support the claim that, on average, people are more than 10 mm taller in the morning than at night.     We reject the claim that, on average, people are more than 10 mm taller in the morning than at night.We have proven that, on average, people are more than 10 mm taller in the morning than at night.

    
    
AM Height (x) PM Height (y) (x - y)
1527 1518 9
1420 1408 12
1439 1432 7
1416 1407 9
1582 1575 7
1457 1444 13
1778 1765 13
1667 1655 12
1676 1664 12
1631 1620 11
1443 1429 14
1403 1392 11
1587 1573 14
1481 1465 16
1752 1742 10
1798 1780 18
1607 1592 15
1604 1585 19
1428 1419 9
1779 1770 9
1581 1572 9
1725 1715 10
1583 1574 9
1703 1696 7
1692 1685 7
1562 1553 9
1643 1631 12
1564 1556 8
1417 1407 10
1592 1579 13
Answer #1

a)  This is a right-tailed test.

b)

SAMPLE 1 SAMPLE 2 difference , Di =sample1-sample2 (Di - Dbar)²
1527 1518 9.000 4.551
1420 1408 12.000 0.751
1439 1432 7.000 17.084
1416 1407 9.000 4.551
1582 1575 7.000 17.084
1457 1444 13.000 3.484
1778 1765 13.000 3.484
1667 1655 12.000 0.751
1676 1664 12.000 0.751
1631 1620 11.000 0.018
1443 1429 14.000 8.218
1403 1392 11.000 0.018
1587 1573 14.000 8.218
1481 1465 16.000 23.684
1752 1742 10.000 1.284
1798 1780 18.000 47.151
1607 1592 15.000 14.951
1604 1585 19.000 61.884
1428 1419 9.000 4.551
1779 1770 9.000 4.551
1581 1572 9.000 4.551
1725 1715 10.000 1.284
1583 1574 9.000 4.551
1703 1696 7.000 17.084
1692 1685 7.000 17.084
1562 1553 9 4.55
1643 1631 12 0.75
1564 1556 8 9.82
1417 1407 10 1.28
1592 1579 13 3.48
sample 1 sample 2 Di (Di - Dbar)²
sum = 47537 47203 334.000 291.467

mean of difference ,    D̅ =ΣDi / n =   11.133                  
                          
std dev of difference , Sd =    √ [ (Di-Dbar)²/(n-1) =    3.1703                  
                          
std error , SE = Sd / √n =    3.1703   / √   30   =   0.5788      
                          
t-statistic = (D̅ - µd)/SE = (   11.13333333   -   10   ) /    0.5788   =   1.96

c)

Degree of freedom, DF=   n - 1 =    29  
  
p-value =        0.0300    [excel function: =t.dist.rt(t-stat,df) ]

d)

p-value>α , fail to reject null hypothesis

e)There is not enough data to support the claim that, on average, people are more than 10 mm taller in the morning than at night.